# Cones of Noether–Lefschetz divisors and moduli spaces of hyperkähler manifolds

### v2

@@ 3.1. Moduli of K3 surfaces
In this case $\Lambda_{2d}=U^{\oplus 2}\oplus E_{8}(-1)^{\oplus 2}\oplus A_{1}(-d)$ and the quotient $\mathcal{F}_{2d}=\mathcal{D}_{\Lambda}/\widetilde{\rm{O}}^{+}\left(\Lambda_{2d}\right)$ is the moduli space of quasi-polarized K3 surfaces of degree $2d$. The method of Section 3 together with the bounds of Theorem 3.7 (and their Sage implementation [heegner_cones]) yield minimal generating rays of ${\rm{Eff}}^{NL}\left(\mathcal{F}_{2d}\right)$ for low $d$. These calculations confirm the predictions of [Pet15]*Remark 4.7.1 and Table 4.5 (aside from one additional generator in the case $d=13$). We record these generators in Table LABEL:sec6:table:NLK3 for $1\leq d\leq 20$.
[…]
@@
The polarization $L$ comes with two invariants singling out a component of the moduli space. These are the Beauville–Bogomolov–Fujiki degree $2d$ and the divisibility $\gamma\in\{1,2\}$. Further, when $\gamma=2$, then $d=4t-1$ for some $t\geq 1$. We will denote by $\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}$ the partial compactification of the corresponding moduli space given by the modular variety $\mathcal{D}_{\Lambda_{h}}\big{/}{\rm{Mon}}^{2}(\Lambda,h)$, where after choosing a marking, $\Lambda_{h}$ is the orthogonal complement of $h=c_{1}(L)$ in $\Lambda$, and ${\rm{Mon}}^{2}(\Lambda,h)=\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$, cf. [Mar11]*Lemma 9.2 and [BBBF23]*Proposition 3.7.
@@
These correspond to the lattice $\Lambda_{h}$ for $(X,L)$ in $\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}$ when $\gamma=1$, respectively $\gamma=2$ with $d=4t-1$. When $\gamma=1$, the discriminant group is isomorphic to $\mathbb{Z}\big{/}2d\mathbb{Z}\times\mathbb{Z}\big{/}2\mathbb{Z}$, generated by $\ell_{*}$ and $\delta_{*}$. When $\gamma=2$, the discriminant group is isomorphic to $\mathbb{Z}/d\mathbb{Z}$ and is generated by $(2u+v)_{*}$.
The Sage program [heegner_cones] following the procedure described above yields the minimal generators of the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}\right)$ in the split case in Table 2 for $d\leq 5$ and in the non-split case in Table 3 for $t\leq 5$ with $d=4t-1$.
@@
[…]
@@ 3.3. Cubic Fourfolds
If $Y\subset\mathbb{P}^{5}$ is a smooth cubic fourfold, then its primitive cohomology $H^{4}(Y,\mathbb{Z})^{\circ}$ together with its intersection form is isomorphic to $\Lambda(-1)=U^{\oplus 2}\oplus E_{8}^{\oplus 2}\oplus A_{2}$ which up to sign, is the same as above with $t=1$. Thus the computation of NL cone of the partial compactification $\mathcal{D}_{\Lambda}/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ (see [Voi86]) of the moduli space of smooth cubic fourfolds has already been computed in Table 3 (with $d=3$ and $\gamma=2$).
@@ 4. Uniruledness
For the orthogonal modular variety $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}(\Lambda)$ consider the degree map
@@
given by taking the degree of a divisor with respect to the Hodge class $\lambda$. The preimage of 00 under this map defines a hyperplane in ${\rm{Pic}}_{\mathbb{Q}}\left(X\right)$. Since all effective divisors have positive degree, the NL cone always lies completely on one side of this hyperplane. In the cases where $K_{X}$ not only lies outside of the NL cone but is in fact negative, meaning that it has negative degree with respect to the Hodge class $\lambda$ and so lies on the other side of this hyperplane, we will conclude the additional statement that the orthogonal modular variety $X$ is uniruled and thus has negative Kodaira dimension. This approach is formalized in Proposition 4.2 below.
@@
$\pi\colon\mathcal{D}_{\Lambda}\rightarrow X=\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right).$
The map $\pi$ is simply ramified [GHS07]*Theorem 2.12 and Corollary 2.13 (see also [GHS13]*Section 6.2) along the union of hyperplanes $D_{\rho}$ such that $\rho\in\Lambda_{\mathbb{Q}}$ is $\widetilde{\rm{O}}^{+}\left(\Lambda\right)$-reflexive, meaning that $\langle\rho,\rho\rangle<0$ and $\sigma_{\rho}$ or $-\sigma_{\rho}$ is in $\widetilde{\rm{O}}^{+}\left(\Lambda\right)$, where $\sigma_{\rho}$ is the reflection given by
@@
Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 3$ splitting off two copies of $U$, let $E_{\frac{n+2}{2},\Lambda}$ be its Eisenstein series, and let ${\rm Br}(\pi)=\sum_{i=1}^{r}\alpha_{m_{i},\mu_{i}}H_{-m_{i},\mu_{i}}$ with $\alpha_{m_{i},\mu_{i}}\in\mathbb{Q}_{>0}$ be the reduced branch divisor of the quotient map $\pi\colon\mathcal{D}_{\Lambda}\rightarrow\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)$. If
@@
then the orthogonal modular variety $X=\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}(\Lambda)$ is uniruled.
@@
from a toroidal compactification $\overline{X}^{\rm tor}$ of $X$ to its Baily–Borel compactification. Note that toroidal compactifications of locally symmetric manifolds of type ${\rm{O}}(2,n)$ are normal and have at worst finite quotient singularities, see [AMRT10]. Since the Hodge class $\lambda$ is ample on $\overline{X}^{\rm BB}$ and $\overline{X}^{\rm BB}\backslash X$ is one-dimensional we can choose a representative for the nef curve class $(\varepsilon^{*}\lambda)^{n-1}$ which does not meet the boundary divisor $\delta$ of the toroidal compactification $\overline{X}^{\rm tor}$.
@@
For any divisor $D$ on $X$, if $\overline{D}$ is the closure of $D$ in $\overline{X}^{\rm tor}$, then since $(\varepsilon^{*}\lambda)^{n-1}$ does not intersect the boundary of $\overline{X}^{\rm tor}$ then the intersection $\overline{D}.(\varepsilon^{*}\lambda)^{n-1}$ is given by the degree of the closure of $D$ in $\overline{X}^{\rm BB}$ with respect to $\lambda$. So consider the map given by the Baily-Borel degree
@@
We denote by $\Lambda_{h}$ the orthogonal complement of $h$ in $\Lambda$. The period domain $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}=\mathcal{D}_{\Lambda_{h}}\big{/}{\rm{O}}^{+}\left(\Lambda,h\right)$ is a partial compactification of the moduli space parameterizing primitively polarized hyperkähler sixfolds of OG6-type with a polarization of degree $2d$ and divisibility $\gamma$. It is always irreducible [Son23]*Section 3 and when $\gamma=1$, is non-empty for all $d\geq 1$ and when $\gamma=2$, is non-empty only for $d\equiv 2,3\mod 4$. Not much is known about the global geometry of the moduli spaces $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$.
In the split case $\gamma=1$, $\Lambda_{h}\cong U^{\oplus 2}\oplus A_{1}(-1)^{\oplus 2}\oplus A_{1}(-d)$. When $\gamma=2$, then $\Lambda_{h}=U^{\oplus 2}\oplus Q_{t}$, where
@@
The polarized monodromy group ${\rm{Mon}}^{2}\left(\Lambda,h\right)\subset{\rm{O}}^{+}\left(\Lambda_{h}\right)$ is given by $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ if $\gamma=2$ and $d=4t-1$, and an index two extension of $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ otherwise. More precisely if $\gamma=1$, or $\gamma=2$ and $d=4t-2$, then
@@
Let $h\in\Lambda$ be an element with $\langle h,h\rangle=2d$ and ${\rm{div}}_{\Lambda}(h)=\gamma$. Since $\Lambda$ and $\Lambda_{h}$ contain two copies of the hyperbolic plane, the map ${\rm{O}}(\Lambda)\longrightarrow{\rm{O}}(D(\Lambda))$ and the respective one for $\Lambda_{h}$ are surjetive. Note that $D(\Lambda)\cong\left(\mathbb{Z}\big{/}2\mathbb{Z}\right)^{\oplus 2}$ generated by $\frac{\delta_{1}}{2}$ and $\frac{\delta_{2}}{2}$. Further, the only non-trivial element in $O(D(\Lambda))$ is the one exchanging the two generators. This in particular implies that $\left[{\rm{O}}^{+}(\Lambda):\widetilde{\rm{O}}^{+}(\Lambda)\right]=2$. Since $\widetilde{\rm{O}}^{+}(\Lambda,h)=\widetilde{\rm{O}}^{+}(\Lambda_{h})$, see for example [ABL24]*Lemma 3.15, either ${\rm{O}}^{+}(\Lambda,h)$ is equal to $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ or it is a double extension.
By Eichler’s Criterion we can always assume $h=e+df$ when $\gamma=1$, and $h=2(e+tf)-\delta$ with $\delta\in\{\delta_{1},\delta_{2},\delta_{1}+\delta_{2}\}$ when $\gamma=2$, the first two happening when $d=4t-1$ and the last one when $d=4t-2$. Note that $\delta_{1}$ and $\delta_{2}$ are in the same ${\rm{Mon}}^{2}(\Lambda)$ orbit so it is enough to consider only one of them. Assume $\gamma=1$ or $\gamma=2$ and $\delta=\delta_{1}+\delta_{2}$. Since $\delta_{1}-\delta_{2}\in\Lambda_{h}$, the reflection $\sigma_{\delta_{1}-\delta_{2}}\in{\rm{O}}^{+}(\Lambda)$ fixes $h$ and exchanges the two generators of $D(\Lambda)$, so $\sigma_{\delta_{1}-\delta_{2}}\in{\rm{O}}^{+}(\Lambda,h)={\rm{Mon}}^{2}\left(\Lambda,h\right)$ and $\sigma_{\delta_{1}-\delta_{2}}\not\in\widetilde{\rm{O}}^{+}(\Lambda_{h})$. Finally, assume $\delta=\delta_{1}$, that is, $\gamma=2$ and $d=4t-1$. If $g\in{\rm{O}}^{+}(\Lambda,h)$, then $g\left(\frac{h}{2}\right)=g(e+tf)-g\left(\frac{\delta_{1}}{2}\right)=\frac{h}{2}$. In particular $g\left(\frac{\delta_{1}}{2}\right)\equiv\frac{\delta_{1}}{2}\mod\Lambda$. This implies $g\in\widetilde{\rm{O}}^{+}\left(\Lambda,h\right)$ and we conclude using the equality induced by restriction $\widetilde{\rm{O}}^{+}\left(\Lambda,h\right)=\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. ∎
@@
    when $\gamma=2$ for $t\leq 10$ and $t=12$ with $d=4t-1$,
@@
    when $\gamma=2$ for $t\leq 9$ and $t=11,13$ with $d=4t-2$.
@@
We will show that in the given cases $X=\mathcal{D}_{\Lambda_{h}}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$ is uniruled. Since $\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)\subset{\rm{Mon}}^{2}(\Lambda,h)$, there is a dominant map $X\longrightarrow\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$ giving us uniruledness for the moduli space $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$. By Proposition 1.2 we just need to verify that $05c_{0,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)+\frac{1}{4}c_{-1,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)<$ 0.
@@
which gives the desired uniruledness of $X$ and hence of $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$. Similarly, when $d=4t-1$ and $\gamma=2$, the highest degree for which we obtain uniruledness is $t=12$. In this case the $E_{0}$ summand of $E_{\frac{7}{2},\Lambda_{h}}$ is
@@
and so $5c_{0,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)+\frac{1}{4}c_{-1,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)=5-\frac{1052352}{4\cdot 51911}<0,$ as desired. In the case $\gamma=2$ and $d=4t-2$ with $t=13$ the $E_{0}$ component of the Eisenstein series of weight $\frac{7}{2}$ corresponding to $\Lambda_{h}$ is

### v3

@@ 3.1. Moduli of K3 surfaces
In this case $\Lambda_{2d}=U^{\oplus 2}\oplus E_{8}(-1)^{\oplus 2}\oplus A_{1}(-d)$ and the quotient $\mathcal{F}_{2d}=\mathcal{D}/\widetilde{\rm{O}}^{+}\left(\Lambda_{2d}\right)$ is the moduli space of quasi-polarized K3 surfaces of degree $2d$. The method of Section 3 together with the bounds of Theorem 3.7 (and their Sage implementation [heegner_cones]) yield minimal generating rays of ${\rm{Eff}}^{NL}\left(\mathcal{F}_{2d}\right)$ for low $d$. These calculations confirm the predictions of [Pet15]*Remark 4.7.1 and Table 4.5 (aside from one additional generator in the case $d=13$). We record these generators in Table LABEL:sec6:table:NLK3 for $1\leq d\leq 20$.
3.2. Hyperkähler fourfolds of ${\rm{K3}}^{[2]}$-type
Let $(X,L)$ be a primitively polarized hyperkähler fourfold of ${\rm{K3}}^{[2]}$-type. The Beauville–Bogomolov–Fujiki lattice $(H^{2}(X,\mathbb{Z}),q_{X})$ is isomorphic to
@@ 3.2. Hyperkähler fourfolds of ${\rm{K3}}^{[2]}$-type
The polarization $L$ comes with two invariants singling out a component of the moduli space. These are the Beauville–Bogomolov–Fujiki degree $2d$ and the divisibility $\gamma\in\{1,2\}$. Further, when $\gamma=2$, then d=4t1d=4t-1 for some $t\geq 1$. We will denote by $\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}$ the partial compactification of the corresponding moduli space given by the modular variety $\mathcal{D}\big/{\rm{Mon}}^{2}(\Lambda,h)$, where after choosing a marking, $\Lambda_{h}$ is the orthogonal complement of $h=c_{1}(L)$ in $\Lambda$, and ${\rm{Mon}}^{2}(\Lambda,h)=\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$, cf. [Mar11]*Lemma 9.2 and [BBBF23]*Proposition 3.7.
@@
These correspond to the lattice $\Lambda_{h}$ for $(X,L)$ in $\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}$ when $\gamma=1$, respectively $\gamma=2$ with d=4t1d=4t-1. When $\gamma=1$, the discriminant group is isomorphic to $\mathbb{Z}\big/2d\mathbb{Z}\times\mathbb{Z}\big/2\mathbb{Z}$, generated by $\ell_{*}$ and $\delta_{*}$. When $\gamma=2$, the discriminant group is isomorphic to $\mathbb{Z}/d\mathbb{Z}$ and is generated by $(2u+v)_{*}$.
The Sage program [heegner_cones] following the procedure described above yields the minimal generators of the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{M}_{{\rm{K3}}^{[2]},2d}^{\gamma}\right)$ in the split case (the case $\gamma=1$) in Table 2 for $d\leq 5$ and in the non-split case (the case $\gamma=2$) in Table 3 for $t\leq 5$ with d=4t1d=4t-1.
@@
Double EPW sextics are particular ramified double covers of certain singular sextic hypersurfaces in $\mathbb{P}^{5}$. When smooth, they are hyperkähler fourfolds of ${\rm{K3}}^{[2]}$-type with polarization of degree $2$ and divisibility $1$ and hence are elements of $\mathcal{M}_{{\rm{K3}}^{[2]},2}^{1}$. Interestingly, the generators $P_{-1,0},P_{-\frac{1}{4},\ell_{*}},P_{-\frac{1}{4},\delta_{*}},P_{-\frac{1}{2},\ell_{*}+\delta_{*}}$ of the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{M}_{{\rm{K3}}^{[2]},2}^{1}\right)$, together with the additional primitive Heegner divisor $P_{-\frac{5}{4},\delta_{*}}$, are precisely the divisorial part of the complement of the image under the period map of the subset of $\mathcal{M}_{{\rm{K3}}^{[2]},2}^{1}$ of double EPW sextics. See [OG15, OG16], [DM19]*Example 6.3.
@@ 3.3. Cubic Fourfolds
If $Y\subset\mathbb{P}^{5}$ is a smooth cubic fourfold, then its primitive cohomology $H^{4}(Y,\mathbb{Z})^{\circ}$ together with its intersection form is isomorphic to $\Lambda(-1)=U^{\oplus 2}\oplus E_{8}^{\oplus 2}\oplus A_{2}$, which, up to sign, is the same as above with $t=1$. Thus the computation of the NL cone of the partial compactification $\mathcal{D}/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$ (see [Voi86]) of the moduli space of smooth cubic fourfolds has been carried out in Table 3 (with $d=3$ and $\gamma=2$).
@@ 4. Uniruledness
For the orthogonal modular variety $X=\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}(\Lambda)$ consider the degree map
@@
given by taking the degree of a divisor with respect to the Hodge class $\lambda$. The preimage of 0 under this map defines a hyperplane in ${\rm{Pic}}_{\mathbb{Q}}\left(X\right)$. Since all effective divisors have positive degree, the NL cone always lies completely on one side of this hyperplane. In the cases where $K_{X}$ not only lies outside of the NL cone but is in fact negative, meaning that it has negative degree with respect to the Hodge class $\lambda$ and so lies on the other side of this hyperplane, we will conclude the additional statement that the orthogonal modular variety $X$ is uniruled and thus has negative Kodaira dimension. This approach is formalized in Proposition 4.2 below.
@@
$\pi\colon\mathcal{D}\rightarrow X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right).$
The map $\pi$ is simply ramified [GHS07]*Theorem 2.12 and Corollary 2.13 (see also [GHS13]*Section 6.2) along the union of hyperplanes $D_{\rho}$ such that $\rho\in\Lambda_{\mathbb{Q}}$ is $\widetilde{\rm{O}}^{+}\left(\Lambda\right)$-reflexive, meaning that $\langle\rho,\rho\rangle<0$ and $\sigma_{\rho}$ or $\sigma_{\rho}$ is in $\widetilde{\rm{O}}^{+}\left(\Lambda\right)$, where $\sigma_{\rho}$ is the reflection given by
@@
Let $\Lambda$ be an even lattice of signature $(2,n)$ with $n\geq 3$ splitting off two copies of $U$, let $E_{\frac{n+2}{2},\Lambda}$ be its Eisenstein series, and let ${\rm Br}(\pi)=\sum_{i=1}^{r}\alpha_{m_{i},\mu_{i}}H_{-m_{i},\mu_{i}}$ with $\alpha_{m_{i},\mu_{i}}\in\mathbb{Q}_{>0}$ be the reduced branch divisor of the quotient map $\pi\colon\mathcal{D}\rightarrow\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)$. If
@@
then the orthogonal modular variety $X=\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}(\Lambda)$ is uniruled.
@@
from a toroidal compactification $\overline{X}^{\rm tor}$ of $X$ to its Baily–Borel compactification. Note that toroidal compactifications of locally symmetric manifolds of type ${\rm{O}}(2,n)$ are normal and one can always choose the toroidal data in such a way that the corresponding cones are simplicial, ensuring that they have at worst finite quotient singularities, see [AMRT10]. Since the Hodge class $\lambda$ is ample on $\overline{X}^{\rm BB}$ and $\X\overline{X}^{\rm BB}\backslash$ X is one-dimensional we can choose a representative for the nef curve class $(\varepsilon^{*}\lambda)^{n-1}$ which does not meet the boundary divisor $\delta$ of the toroidal compactification $\overline{X}^{\rm tor}$.
@@
For any divisor $D$ on $X$, if $\overline{D}$ is the closure of $D$ in $\overline{X}^{\rm tor}$, since $(\varepsilon^{*}\lambda)^{n-1}$ does not intersect the boundary of $\overline{X}^{\rm tor}$, then the intersection $\overline{D}.(\varepsilon^{*}\lambda)^{n-1}$ is given by the degree of the closure of $D$ in $\overline{X}^{\rm BB}$ with respect to $\lambda$. So consider the map given by the Baily-Borel degree
@@ 4.1. Moduli of $\rm{OG6}$}-type hyperkähler manifolds
We denote by $\Lambda_{h}$ the orthogonal complement of $h$ in $\Lambda$. The period domain $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}=\mathcal{D}\big/{\rm{Mon}}^{2}\left(\Lambda,h\right)$ is a partial compactification of the moduli space parameterizing primitively polarized hyperkähler sixfolds of OG6-type with a polarization of degree $2d$ and divisibility $\gamma$. It is always irreducible [Son23]*Section 3. Moreover, when $\gamma=1$ it is non-empty for all $d\geq 1$ and when $\gamma=2$ it is non-empty only for $d\equiv 2,3\mod 4$. Not much is known about the global geometry of the moduli spaces $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$.
In the split case $\gamma=1$, we have $\Lambda_{h}\cong U^{\oplus 2}\oplus A_{1}(-1)^{\oplus 2}\oplus A_{1}(-d)$. When $\gamma=2$, then $\Lambda_{h}=U^{\oplus 2}\oplus Q_{t}$, where
@@
The polarized monodromy group ${\rm{Mon}}^{2}\left(\Lambda,h\right)\subset{\rm{O}}^{+}\left(\Lambda_{h}\right)$ is given by $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ if $\gamma=2$ and d=4t1d=4t-1, and an index two extension of $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ otherwise. More precisely if $\gamma=1$, or $\gamma=2$ and d=4t2d=4t-2, then
@@
Let $h\in\Lambda$ be an element with $\langle h,h\rangle=2d$ and ${\rm{div}}_{\Lambda}(h)=\gamma$. Since $\Lambda$ and $\Lambda_{h}$ contain two copies of the hyperbolic plane, the map ${\rm{O}}(\Lambda)\longrightarrow{\rm{O}}(D(\Lambda))$ and the respective one for $\Lambda_{h}$ are surjective. Note that $D(\Lambda)\cong\left(\mathbb{Z}\big/2\mathbb{Z}\right)^{\oplus 2}$, generated by $\frac{\delta_{1}}{2}$ and $\frac{\delta_{2}}{2}$. Further, the only non-trivial element in ${\rm O}(D(\Lambda))$ is the one exchanging the two generators. This in particular implies that $\left[{\rm{O}}^{+}(\Lambda):\widetilde{\rm{O}}^{+}(\Lambda)\right]=2$. Since $\widetilde{\rm{O}}^{+}(\Lambda,h)=\widetilde{\rm{O}}^{+}(\Lambda_{h})$, see for example [ABL24]*Lemma 3.15, either ${\rm{O}}^{+}(\Lambda,h)$ is equal to $\widetilde{\rm{O}}^{+}(\Lambda_{h})$ or it is a double extension.
By Eichler’s Criterion we can always assume $h=e+df$ when $\gamma=1$, and $h=2(e+tf)-\delta$ with $\delta\in\{\delta_{1},\delta_{2},\delta_{1}+\delta_{2}\}$ when $\gamma=2$, the first two happening when d=4t1d=4t-1 and the last one when d=4t2d=4t-2. Note that $\delta_{1}$ and $\delta_{2}$ are in the same ${\rm{Mon}}^{2}(\Lambda)$ orbit so it is enough to consider only one of them. Assume $\gamma=1$ or $\gamma=2$ and $\delta=\delta_{1}+\delta_{2}$. Since $\delta_{1}-\delta_{2}\in\Lambda_{h}$, the reflection $\sigma_{\delta_{1}-\delta_{2}}\in{\rm{O}}^{+}(\Lambda)$ fixes $h$ and exchanges the two generators of $D(\Lambda)$, so $\sigma_{\delta_{1}-\delta_{2}}\in{\rm{O}}^{+}(\Lambda,h)={\rm{Mon}}^{2}\left(\Lambda,h\right)$ and $\sigma_{\delta_{1}-\delta_{2}}\not\in\widetilde{\rm{O}}^{+}(\Lambda_{h})$. Finally, assume $\delta=\delta_{1}$, that is, $\gamma=2$ and d=4t1d=4t-1. If $g\in{\rm{O}}^{+}(\Lambda,h)$, then $g\left(\frac{h}{2}\right)=g(e+tf)-g\left(\frac{\delta_{1}}{2}\right)=\frac{h}{2}$. In particular $g\left(\frac{\delta_{1}}{2}\right)\equiv\frac{\delta_{1}}{2}\mod\Lambda$. This implies $g\in\widetilde{\rm{O}}^{+}\left(\Lambda,h\right)$ and we conclude using the equality induced by restriction $\widetilde{\rm{O}}^{+}\left(\Lambda,h\right)=\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$. ∎
@@
    when $\gamma=2$ for $t\leq 10$ and $t=12$ with d=4t1d=4t-1,
@@
    when $\gamma=2$ for $t\leq 9$ and $t=11,13$ with d=4t2d=4t-2.
@@
We will show that in the given cases $X=\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)$ is uniruled. Since $\widetilde{\rm{O}}^{+}\left(\Lambda_{h}\right)\subset{\rm{Mon}}^{2}(\Lambda,h)$, there is a dominant map $X\longrightarrow\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$ giving us uniruledness for the moduli space $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$. By Proposition 1.2 we just need to verify that $5c_{0,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)+\frac{1}{4}c_{-1,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)<0$.
@@
which gives the desired uniruledness of $X$ and hence of $\mathcal{M}_{{\rm{OG6}},2d}^{\gamma}$. Similarly, when d=4t1d=4t-1 and $\gamma=2$, the highest degree for which we obtain uniruledness is $t=12$. In this case the $E_{0}$ summand of $E_{\frac{7}{2},\Lambda_{h}}$ is
@@
and so $5c_{0,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)+\frac{1}{4}c_{-1,0}\left(E_{\frac{7}{2},\Lambda_{h}}\right)=5-\frac{1052352}{4\cdot 51911}<0,$ as desired. In the case $\gamma=2$ and d=4t2d=4t-2 with $t=13$ the $E_{0}$ component of the Eisenstein series of weight $\frac{7}{2}$ corresponding to $\Lambda_{h}$ is